OpenStudy (anonymous):

Please Help!! Find the equation of the tangent line to the following curve at the point (0, pi) sin(x+y)=2x

OpenStudy (anonymous):

do you know about inverse trig functions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i've tried everything can you please get me the answer?

OpenStudy (anonymous):

@RadEn just helped me hopefully he can help you . . .otherwise I'll let you know if I have an idea

OpenStudy (anonymous):

ok

OpenStudy (raden):

we have to derive implicitly !

OpenStudy (anonymous):

how do i do that?

OpenStudy (raden):

sin(x+y)=2x cos(x+y) + cos(x+y) y' = 2 cos(x+y) (1 + y') = 2 1 + y ' = 2/ cos(x+y) y ' = [2/cos(x+y)] - 1

OpenStudy (raden):

now, subtitute x=0 and y=pi to equation above to get the slope (m)

OpenStudy (anonymous):

y-y1=m(x-x1)?

OpenStudy (raden):

m = y '(0,pi) = [2/cos(0+pi)] - 1 = 2/cospi - 1 = 2/(-1) - 1 = -2-1=-3

OpenStudy (anonymous):

got it!!!!!!

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

y=-3x+pi

OpenStudy (raden):

it just a slope to get the slope equation use the formula like u said : y-y1=m(x-x1) EXACTLY, U ARE RIGHT :)

OpenStudy (anonymous):

can i ask you one more question?

OpenStudy (anonymous):

yea sorry about that I knew it was either inverse trig functions or implicit differentiation but I could not remember how to do implicit at the moment

OpenStudy (raden):

sure, but close this, and posting in the new question :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I would have given that medal to @RAdEn he is the one who helped you

OpenStudy (anonymous):

thank you !!